Derive the one-dimensional wave
equation.
1
Expert's answer
2021-09-07T03:08:53-0400

Consider the relation between Newton’s law that is applied to the volume "\Delta V" in the direction "x" :

"\Delta F=\Delta m\frac{dv_{x}}{dt}" (Newton' law)

Where, "F" : force acting on the element with volume "\Delta V"

"\Delta F_{x}=-\Delta p_{x}\Delta S_{x}"

"=(\frac{\partial p}{\partial x}\Delta x+\frac{\partial p}{\partial x}dt)\Delta S_{x}"

"\simeq -\frac{\partial p}{\partial x}\Delta V-\Delta V\frac{\partial p}{\partial x}=\Delta m\frac{dv_{x}}{dt}"

as dt is small, it is not considered and "\Delta Sx" is in "x" direction so ΔyΔz and from Newton’s law

"=\rho \Delta V\frac{dv_{x}}{dt}"

From, "\frac{dv_{x}}{dt}\ \text{as}\ \frac{\partial v_{x}}{\partial t}"

"\frac{dv_{x}}{dt}=\frac{\partial v_{x}}{\partial t}+v_{x}\frac{\partial v_{x}}{\partial x}\approx \frac{\partial v_{x}}{\partial x}-\frac{\partial p}{\partial x}=\rho \frac{\partial v_{x}}{\partial t}"

Above equation is known as equation of motion.

"-\frac{\partial }{\partial x}(\frac{\partial p}{\partial x})=\frac{\partial }{\partial x}(\rho \frac{\partial v_{x}}{\partial t})"

"-\frac{\partial^2 p}{\partial x^2}=\rho \frac{\partial }{\partial t}(-\frac{1}{K}\frac{\partial p}{\partial t})" (From conservation of mass)

"\frac{\partial p^{2}}{\partial x^{2}}-\frac{\rho }{K}\frac{\partial^2 p}{\partial t^2}=0"

Where, K: bulk modulus

Rewriting the above equation:

"\frac{\partial p^{2}}{\partial x^{2}}-\frac{1}{c}^{2}\frac{\partial^2 p}{\partial t^2}=0"

Where, c: velocity of sound given as "c=\sqrt{\frac{K}{\rho }}"

Thus, above is the one-dimensional wave equation derivation.